4 ms·
1. Bet $10000 on black at the casino. If you win, well done. If you lose continue. 2. Bet $20000 on black at the casino, if you win, well done. If you lose, yo
by meric 9y ago
1. Bet $10000 on black at the casino. If you win, well done. If you lose continue.
2. Bet $20000 on black at the casino, if you win, well done. If you lose, you lost $30000.
There is 75% chance to make $10000 - casino edge using this strategy. You can do this on any day of the week the casino is open.
- amingilani 9y agoDoubling down, also called a martingale strategy. This was awesome while it lasted. I used this while playing with dogecoin, the trouble started when I automated it using a script. Trouble is, even when the chance of losing is 0.25%, you'll lost once every 400 games. And at one point, the loss will be greater than all the money you had.. and there's no coming back from that.
- csomar 9y agoHe said 75% chance, so odds are against you.
- amingilani 9y agoActually, he meant a 75% chance of winning.. Given the chance of winning a round is p=0.5, the chance of losing is q = 1 - p = 0.5 If you've lost the first round, the odds of you twice in a row are q = 0.5^(2 rounds) = 0.25 Which means the odds of you winning are p = 1 - q = 1 - 0.25 = 0.75.
- 0x8BADF00D 9y agoRight, but the expected value of the gamble is negative.
- dfinch 9y agoExpected value is 0, right? .75 * 10000 - .25 * 30000
- amingilani 9y agoDoubling down, also called a martingale strategy. This was awesome while it lasted. I used this while playing with dogecoin, the trouble started when I automated it using a script. Trouble is, even when the chance of losing is 0.25%, you'll lost once every 400 games. And at one point, the loss will be greater than all the money you had.. and there's no coming back from that.
- stadeschuldt 9y agoThis won't work: https://blog.tafkas.net/2009/06/09/a-not-so-safe-betting-strategy-for-winning-at-roulette/ https://blog.tafkas.net/2009/06/09/a-not-so-safe-betting-str...